OpenStudy (anonymous):

Condense . express as a single logarithm and if possible ,simplify logb^8x+6(logb^x-logb^y)

sam (.sam.):

What do you get? Try using these rules http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (anonymous):

i don't know how to start

sam (.sam.):

Is it \[\large \log b^{8x}+6(\log b^x-\log b^y)\]

OpenStudy (anonymous):

right

sam (.sam.):

Distribute 6 first

OpenStudy (anonymous):

Give me more hint plzz

sam (.sam.):

Use \[\log a^b + \log a^d = \log (a^ba^d)=\log(a^{b+d})\] \[\log a^b -\log a ^d=\log(\frac{a^b}{a^d})=\log(a^{b-d})\]